# On Speeding Up Language Model Evaluation

### v1

@@ 6 Discussion and Conclusion
One limitation of our proposed algorithms is that we assume there is a two-dimensional fixed-size matrix as our dataset for active method-example selection. In some real-world applications, new rows or columns can be gradually incorporated and the matrix size is dynamic. In other scenarios such as Chatbot Arena, instead of being able to decide what examples to select, we can only select a pair of models to compare for a user-specified example. We leave studying how to minimize budget to find the best method for these new settings as future work. Other promising extensions of our work are to apply more advanced best-arm identification algorithm in the literature, and fixed confidence setting instead of fixed budget setting.
In conclusion, we formulate the problem that when the evaluation for each model-example is resource intensive in regard to money, compute and time, given a certain budget of resource, how can we have still have a great chance to identify the best model on a specific dataset among a set of candidates. We propose two algorithms to tackle the problem, which all sequentially decide which model-example pair to evaluate next by observing the previous evaluated pairs. The first algorithm follows the idea of a classic multi-arm bandit algorithm UCB-E and enjoys the similar theoretical guarantee to lower bound the chance given any budget. The second algorithm, UCB-E-LRF extends UCB-E by leveraging the finding of approximate low-rank of the target evaluation matrix. In the experiment, we show that both algorithms introduced in our paper are significantly better than one that just uniformly samples method-example pairs. Moreover, we identified the condition when the UCB-E or UCB-E-LRF works better than the other. We do not foresee any negative societal impacts for our work.
@@ 7 Acknowledgement
We thank Justin Lovelace and Varsha Kishore for their helpful discussion and feedback on the paper draft. JPZ is supported by grant from the Natural Sciences and Engineering Research Council of Canada (NSERC) (567916). CKB and CPG are support by the Cornell University AI for Science Institute, the AI Climate Institute, the National Science Foundation, the National Institute of Food and Agriculture, and the Air Force Office of Scientific Research. RW is supported by grants from National Science Foundation NSF (CIF-2402817, CNS-1804829), SaTC-2241100, CCF-2217058, ARO-MURI (W911NF2110317), and ONR under N00014-24-1-2304. Wen Sun is supported by NSF IIS-2154711 and NSF CAREER 2339395. This research is also supported by grants from the National Science Foundation NSF (IIS-2107161, and IIS-1724282, HDR-2118310), the Cornell Center for Materials Research with funding from the NSF MRSEC program (DMR-1719875), DARPA, arXiv, LinkedIn, and the New York Presbyterian Hospital.
Appendix A Appendix
A.1 Additional Description of Datasets Used
Figure 5: Histogram of model performance on all six datasets along with their $H_{1}$ values. It can be seen that datasets such as AlpacaEval have larger gap between the best and second best method leads to a much smaller $H_{1}$, making identifying the best method easier.
@@ A.1 Additional Description of Datasets Used
To simulate a prompt engineering use case, we create two datasets: GSM8K Prompts and PIQA Prompts. For these datasets, we ask GPT4 to generate 205 and 177 prompts following the prompt engineering work from Yang et al. [60]. The LLM used to perform inference on the datasets are Mistra-7B [32] and Tulu-7B [58] respectively. We evaluate the prompts on 784 and 1546 questions on the training set (about 10% of the size of the two training sets). For scoring, we extract the final answers from the responses generated by the LLMs and compared them with the ground truth answers (for GSM8K) or ground truth choice (for PIQA).
@@
A.2 Algorithmic Description of Baselines and Ablations
@@
Input: The evaluation budget $T$, a set of methods $\mathcal{F}$ and a set of examples $\mathcal{X}$.
@@
1:  $\forall i$, the set of evaluated examples $S_{i}=\{\}$.
@@
3:     Select: Draw $i_{t}\in[n]$; Draw $j_{t}\in[m]\backslash S_{i}$.
4:     Evaluate: Run inference for the method-example pair $(f_{i_{t}},x_{j_{t}})$, send the result to the annotator, and receive $e(f_{i_{t}},x_{j_{t}})$; $E^{\rm obs}_{i_{t},j_{t}}\leftarrow e(f_{i_{t}},x_{j_{t}})$.
5:     Update: $S_{i_{t}}\leftarrow S_{i_{t}}\cup\{j_{t}\}$.
@@
Input: The evaluation budget $T$, a set of methods $\mathcal{F}$ and a set of examples $\mathcal{X}$.
@@
1:  the set of evaluated examples (same for all methods) $S=\{\}$.
2:  for $t=1,\cdots,[T/n]$ do
3:     Select: $I_{t}=\{1,\cdots,n\}$; Draw $j_{t}\in[m]\backslash S$.
4:     for $i_{t}\in I_{t}$ do
5:        Evaluate: Run inference for the method-example pair $(f_{i_{t}},x_{j_{t}})$, send the result to the annotator, and receive $e(f_{i_{t}},x_{j_{t}})$; $E^{\rm obs}_{i_{t},j_{t}}\leftarrow e(f_{i_{t}},x_{j_{t}})$
@@
7:     Update: $S_{t}\leftarrow S_{t}\cup\{j_{t}\}$.
8:  end for
Return: $\hat{i}^{*}=\arg\max_{i}\frac{\sum_{j\in S}E^{\rm obs}_{i,j}}{|S|}$
Algorithm 5 LRF ($\mathcal{A}_{lrf}(T,\mathcal{F},\mathcal{X};\mathcal{M},r,K,T_{0},b)$)
Input: The evaluation budget $T$, a set of methods $\mathcal{F}$ and a set of examples $\mathcal{X}$, the LRF oracle $\mathcal{M}$, the rank $r$, # solutions in ensemble $K$, the warm-up budget $T_{0}$, the batch size $b$.
@@
1:  Uniformly sample $T_{0}$ method-example pairs from $[n]\times[m]$ and get the observed scoring matrix $E^{\rm obs}\in\left([0,1]\cup\{?\}\right)^{n\times m}$ and the observation matrix $O\in\{0,1\}^{n\times m}$ w.r.t. these $T_{0}$ evaluations.
2:  for $t=T_{0},\cdots,T_{0}+[(T-T_{0})/b]$ do
3:     Select: Draw $I_{t}$ by sampling $b$ distinct indices uniformly at random from $[n]$; Draw $J_{t}$ by sampling $b$ distinct indices uniformly at random from $[m]$.
4:     for $(i_{t},j_{t})\in(I_{t},J_{t})$ do
5:        Evaluate: Run inference for the method-example pair $(f_{i_{t}},x_{j_{t}})$, send the result to the annotator, and receive $e(f_{i_{t}},x_{j_{t}})$; $E^{\rm obs}_{i_{t},j_{t}}\leftarrow e(f_{i_{t}},x_{j_{t}})$
@@
7:     Update: $\forall j_{t}\in J_{t}$, $O_{i_{t},j_{t}}\leftarrow 1$; $\hat{E}\leftarrow\mathcal{M}(E^{\rm obs},O;r,K)$.
8:  end for
Return: $\hat{i}^{*}=\arg\max_{i}\frac{1}{m}\sum_{j=1}^{m}(O_{i,j}E^{\rm obs}_{i,j}+(1-O_{i,j})\hat{E}_{i,j})$
Algorithm 6 UCB-E-LRF (Score Only) ($\mathcal{A}_{uel-so}(T,\mathcal{F},\mathcal{X};\mathcal{M},r,K,T_{0},b,\eta)$)
Input: The evaluation budget $T$, a set of methods $\mathcal{F}$ and a set of examples $\mathcal{X}$, the LRF oracle $\mathcal{M}$, the rank $r$, # solutions in ensemble $K$, the warm-up budget $T_{0}$, the batch size $b$, the uncertainty scaling $\eta$.
@@
1:  Uniformly sample $T_{0}$ method-example pairs from $[n]\times[m]$ and get the observed scoring matrix $E^{\rm obs}\in\left([0,1]\cup\{?\}\right)^{n\times m}$ and the observation matrix $O\in\{0,1\}^{n\times m}$ w.r.t. these $T_{0}$ evaluations.
2:  $\hat{E}\leftarrow\mathcal{M}(E^{\rm obs},O)$, $\forall i,B_{i}\leftarrow\frac{1}{m}\sum_{j=1}^{m}(O_{i,j}E^{\rm obs}_{i,j}+(1-O_{i,j})\hat{E}_{i,j})+\sqrt{\frac{\eta}{\sum_{j=1}^{m}(O_{i,j})}}$
3:  for $t=T_{0},\cdots,T_{0}+[(T-T_{0})/b]$ do
4:     Select: Draw $i_{t}\in\arg\max_{i}B_{i}$; Draw $J_{t}$ by sampling $b$ distinct indices uniformly at random from $[m]$.
5:     for $j_{t}\in J_{t}$ do
6:        Evaluate: Run inference for the method-example pair $(f_{i_{t}},x_{j_{t}})$, send the result to the annotator, and receive $e(f_{i_{t}},x_{j_{t}})$; $E^{\rm obs}_{i_{t},j_{t}}\leftarrow e(f_{i_{t}},x_{j_{t}})$
@@
8:     Update: $\forall j_{t}\in J_{t}$, $O_{i_{t},j_{t}}\leftarrow 1$; $\hat{E}\leftarrow\mathcal{M}(E^{\rm obs},O;r,K)$; $\forall f_{i}\in\mathcal{F},B_{i}\leftarrow\frac{1}{m}\sum_{j=1}^{m}(O_{i,j}E^{\rm obs}_{i,j}+(1-O_{i,j})\hat{E}_{i,j})+\sqrt{\frac{\eta}{\sum_{j=1}^{m}(O_{i,j})}}$.
9:  end for
Return: $\hat{i}^{*}=\arg\max_{i}\frac{1}{m}\sum_{j=1}^{m}(O_{i,j}E^{\rm obs}_{i,j}+(1-O_{i,j})\hat{E}_{i,j})$

### v2

@@ 5 Related Work
LLM evaluation with multi-arm bandits.
Concurrently, Shi et al. [56] explores using multi-arm bandit for prompt optimization in a limited budget setting. Our work does not solely focus on prompt optimization and instead a more general evaluation setup. We also propose an algorithm that leverages low-rank factorization for better best-arm identification performance.
@@ 6 Discussion and Conclusion
One limitation of our proposed algorithms is that we assume there is a two-dimensional fixed-size matrix as our dataset for active method-example selection. In some real-world applications, new rows or columns can be gradually incorporated and the matrix size is dynamic. In other scenarios such as Chatbot Arena, instead of being able to decide what examples to select, we can only select a pair of models to compare for a user-specified example. We leave studying how to minimize budget to find the best method for these new settings as future work.
In this work we focus on identifying the top-performer with limited budget as it is a very common problem; covering a wide range of applications particularly for practitioners who simply want to find the best prompt or model for deployment purposes. Another interesting direction for future work is recovering a full or top K ranking of performers, as it provides more fine-grained information that is necessary for specific use cases such as iterating on LLM design, and tracking LLM ranking over time. Of course in this setting the opportunity for resource savings is less than in the top-performer setting. In Figure 3, we evaluate our algorithms on NDCG@K as preliminary evidence supporting the possibility of resource saving recovering a top K ranking. We leave a more thorough study for future work.
In conclusion, we formulate the problem that when the evaluation for each model-example is resource intensive in regard to money, compute and time, given a certain budget of resource, how can we still have a high probability of identifying the best model on a specific dataset among a set of candidates. We propose two algorithms to tackle the problem, which all sequentially decide which model-example pair to evaluate next by observing the previous evaluated pairs. The first algorithm follows the idea of a classic multi-arm bandit algorithm UCB-E and enjoys the similar theoretical guarantee to lower bound the chance given any budget. The second algorithm, UCB-E-LRF extends UCB-E by leveraging the finding of approximate low-rank of the target evaluation matrix. In the experiment, we show that both algorithms introduced in our paper are significantly better than one that just uniformly samples method-example pairs. Moreover, we identified the condition when the UCB-E or UCB-E-LRF works better than the other. We do not foresee any negative societal impacts for our work.
@@ 7 Acknowledgement
We thank Justin Lovelace, Varsha Kishore, and Katie Luo for their helpful discussion and feedback on the paper draft. JPZ is supported by a grant from the Natural Sciences and Engineering Research Council of Canada (NSERC) (567916). CKB and CPG are supported by the Cornell University AI for Science Institute, the AI Climate Institute, the National Science Foundation, the National Institute of Food and Agriculture, and the Air Force Office of Scientific Research. RW is supported by grants from National Science Foundation NSF (CIF-2402817, CNS-1804829), SaTC-2241100, CCF-2217058, ARO-MURI (W911NF2110317), and ONR under N00014-24-1-2304. Wen Sun is supported by NSF IIS-2154711 and NSF CAREER 2339395. This research is also supported by grants from the National Science Foundation NSF (IIS-2107161, and IIS-1724282, HDR-2118310), the Cornell Center for Materials Research with funding from the NSF MRSEC program (DMR-1719875), DARPA, arXiv, LinkedIn, and the New York Presbyterian Hospital.
Appendix A Additional Details - Datasets
Figure 5: Bar plot of model performance on all examples in all six datasets. Examples are ordered by their rank. Dataset $H_{1}$ values are displayed next to dataset name. It can be seen that datasets, such as AlpacaEval, that have a large gap between the best and second best method have small $H_{1}$ values. The $H1$ value indicates the difficulty of identifying the best method with smaller values indicating easier identification.
@@ Appendix A Additional Details - Datasets
To simulate a prompt engineering use case, we create two datasets: GSM8K Prompts and PIQA Prompts. For these datasets, we ask GPT4 to generate 205 and 177 prompts following the prompt engineering work from Yang et al. [60]. The LLM used to perform inference on the datasets are Mistra-7B [32] and Tulu-7B [59] respectively. We evaluate the prompts on 784 and 1546 questions on the training set (about 10% of the size of the two training sets). For scoring, we extract the final answers from the responses generated by the LLMs and compare them with the ground truth answers (for GSM8K) or ground truth choice (for PIQA).
@@
Table 3: Ratio of the largest to second largest eigenvalue.
Dataset Name GSM8K Prompts PIQA Prompts GSM8K Models PIQA Models AlpacaEval (Drop Annotator) AlpacaEval
$\frac{\lambda_{1}(E)}{\lambda_{2}(E)}$ 6.0729 11.2547 3.0048 5.0712 2.8573 2.7316
Figure 6: Bar plot of the explained variance of each principal component for all six datasets.
Appendix B Additional Details - Algorithms
@@
Input: The evaluation budget $T$, a set of methods $\mathcal{F}$, a set of examples $\mathcal{X}$.
@@
1:  The observation matrix $O:=\{0\}^{n\times m}$, the observed scoring matrix $E^{\rm obs}:=\{?\}^{n\times m}$.
@@
3:     Select: Draw uniformly at random $(i,j)\in\{[n]\times[m]|O_{ij}=0\}$.
4:     Evaluate: Run inference for the method-example pair $(f_{i},x_{j})$, score the result, and receive $e(f_{i}(x_{j}))$; $E^{\rm obs}_{ij}\leftarrow e(f_{i}(x_{j}))$.
5:     Update: $O_{i,j}\leftarrow 1$;
@@
Input: The evaluation budget $T$, a set of methods $\mathcal{F}$, a set of examples $\mathcal{X}$.
@@
1:  The observation matrix $O:=\{0\}^{n\times m}$, the observed scoring matrix $E^{\rm obs}:=\{?\}^{n\times m}$, initial random example index $j$.
2:  for $t=1,\cdots,T$ do
3:     Select: Draw uniformly at random $i\in\{[n]|O_{ij}=0\}$.
4:     Evaluate: Run inference for the method-example pair $(f_{i},x_{j})$, score the result, and receive $e(f_{i}(x_{j}))$; $E^{\rm obs}_{ij}\leftarrow e(f_{i}(x_{j}))$.
5:     Update: $O_{i,j}\leftarrow 1$; If $\sum_{i=1}^{n}O_{ij}=n$, draw uniformly at random $j\in\{[m]|\sum_{i=1}^{n}O_{ij}=0\}$.
@@
Return: $\hat{i}^{*}=\arg\max_{i}\frac{\sum_{j=1}^{m}O_{ij}\cdot E^{\rm obs}_{ij}}{\sum_{j=1}^{m}O_{ij}}.$
Algorithm 5 LRF ($\mathcal{A}_{lrf}(T,\mathcal{F},\mathcal{X};\mathcal{M},r,S,T_{0})$)
Input: The evaluation budget $T$, a set of methods $\mathcal{F}$, a set of examples $\mathcal{X}$, the low-rank factorization model $\mathcal{M}$, the rank $r$, the ensemble size $S$, the warm-up budget $T_{0}$.
@@
1:  Uniformly draw $T_{0}$ method-example pairs from $[n]\times[m]$ and get the observation matrix $O\in\{0,1\}^{n\times m}$ and observed scoring matrix $E^{\rm obs}\in\left([0,1]\cup\{?\}\right)^{n\times m}$ w.r.t. these $T_{0}$ evaluations.
2:  for $t=T_{0},\cdots,T$ do
3:     Select: Draw uniformly at random $(i,j)\in\{[n]\times[m]|O_{ij}=0\}$.
4:     Evaluate: Run inference for the method-example pair $(f_{i},x_{j})$, score the result, and receive $e(f_{i}(x_{j}))$; $E^{\rm obs}_{ij}\leftarrow e(f_{i}(x_{j}))$.
5:     Update: $O_{i,j}\leftarrow 1$; $\hat{E}\leftarrow\mathcal{M}(E^{\rm obs},O;r,S)$.
@@
Return: $\hat{i}^{*}=\arg\max_{i}\frac{1}{m}\sum_{j=1}^{m}(O_{ij}E^{\rm obs}_{ij}+(1-O_{ij})\hat{E}_{ij}).$
Algorithm 6 UCB-E-LRF (Score Only) ($\mathcal{A}_{uel-so}(T,\mathcal{F},\mathcal{X};\mathcal{M},r,S,T_{0},\eta)$)
Input: The evaluation budget $T$, a set of methods $\mathcal{F}$, a set of examples $\mathcal{X}$, the low-rank factorization model $\mathcal{M}$, the rank $r$, the ensemble size $S$, the warm-up budget $T_{0}$, the uncertainty scaling $\eta$.
@@
1:  Uniformly draw $T_{0}$ method-example pairs from $[n]\times[m]$ and get the observation matrix $O\in\{0,1\}^{n\times m}$ and observed scoring matrix $E^{\rm obs}\in\left([0,1]\cup\{?\}\right)^{n\times m}$ w.r.t. these $T_{0}$ evaluations.
2:  $\hat{E},R\leftarrow\mathcal{M}(E^{\rm obs},O;r,S)$; $\forall f_{i}\in\mathcal{F},B_{i}\leftarrow\frac{1}{m}\sum_{j=1}^{m}(O_{ij}E^{\rm obs}_{ij}+(1-O_{ij})\hat{E}_{ij}+\eta R_{ij})$
3:  for $t=T_{0},\cdots,T$ do
4:     Select: Draw uniformly at random $i\in\arg\max_{k\,\mid\,(\sum_{j=1}^{m}O_{kj})\neq m}B_{k}$; Draw uniformly at random $j\in\{k\in[m]|O_{ik}=0\}$.
5:     Evaluate: Run inference for the method-example pair $(f_{i},x_{j})$, score the result, and receive $e(f_{i}(x_{j}))$; $E^{\rm obs}_{ij}\leftarrow e(f_{i}(x_{j}))$.
6:     Update: $O_{i,j}\leftarrow 1$; $\hat{E}\leftarrow\mathcal{M}(E^{\rm obs},O;r,S)$; $\forall f_{i}\in\mathcal{F},B_{i}\leftarrow\frac{1}{m}\sum_{j=1}^{m}(O_{ij}E^{\rm obs}_{ij}+(1-O_{ij})\hat{E}_{ij}+\eta R_{ij})$.
@@
Return: $\hat{i}^{*}=\arg\max_{i}\frac{1}{m}\sum_{j=1}^{m}(O_{ij}E^{\rm obs}_{ij}+(1-O_{ij})\hat{E}_{ij}).$